OpenStudy (anonymous):

Another question... rearrange into quadratic form ax^2+bx+c=0 0.20= x^2/(25-x) where a=1, and identify values of b and c

OpenStudy (anonymous):

multiply both side by (25-x), and move all the terms to the desired side. coefficient before x^2 is a, coefficient before x is b, and constant term is c.

OpenStudy (anonymous):

then you need to divide the whole eqn by a number to yield a= 1. so our final b and final c is divided from the original b and c's by that amount.

OpenStudy (anonymous):

Got to here then I got confused: 0.20=x^2/ (25-x) 0.20 (25-x)=x^2 5- 0.20x=x^2

OpenStudy (anonymous):

lets have all the babies on our right side of the equal sign. moving everything to the right side, you get x^2+0.20x-5=0

OpenStudy (anonymous):

Shouldn't your 0.20x be negative?

OpenStudy (anonymous):

Because you distribute the 0.20 to (25-x) So wouldn't it read x^2-0.20+5?

OpenStudy (jhannybean):

\[\large \color{red}ax^2 +\color{red}bx+\color{red}c=0\]\[\large 0.20 = \frac{x^2}{25-x}\]\[\large 0.20(25-x) = x^2 \]\[\large 0.20(25) - 0.20x = x^2 \]\[\large \color{red}{1}x^2 + \color{red}{0.20}x -\color{red}{0.20(25)}=0\]

OpenStudy (jhannybean):

It's not negative because you're moving it over to the other side of the equal sign.

OpenStudy (anonymous):

but you have: \[0.20(25)-0.20x\] on one line then: \[+0.20x-0.20(25)\]

OpenStudy (jhannybean):

\[\large 0.20(25) - 0.20x = x^2\] get everything to the right side. \[\large 0.2(25) -(0.20(25)) -0.20x =x^2\]\[\large -0.20x = x^2 -0.20(25)\]\[\large -0.20x + 0.20x = x^2 -0.20(25) +0.20x \]\[\large 0 = x^2 +0.20x -0.20(25)\]

OpenStudy (jhannybean):

Understand it now?

OpenStudy (anonymous):

I'm still very confused. Thank you for trying to help me, though. I'm a little fuzzy because I had surgery today. I guess I'll just skip over it and try coming back another time. Thank you again

OpenStudy (jhannybean):

What part are you confused with? What we have to do is try making the left side 0 by subtracting each term by its opposite.

OpenStudy (anonymous):

how you got from \[0.20(25)-0.20x=x^2\] to \[0.20(25)-(0.20(25))-0.20x=x2\]

OpenStudy (jhannybean):

I want to eliminate the "0.20(25)" from the left side by SUBTRACTING it from the left. Whatever i do to the LEFT side i have to do to the RIGHT side.

OpenStudy (anonymous):

It's already on the left.. Why is it being eliminated?

OpenStudy (anonymous):

So now were moving everything to the right?

OpenStudy (jhannybean):

it's being eliminated from the left to the right because we are trying to fit in the format of \(\large ax^2 +bx +c = 0\) in which case our 0 will be on the left and all the other numbers will be on the right

OpenStudy (jhannybean):

Thats why we've written our solutions as \(\large 0 = x^2 +0.20x -0.20(25)\)

OpenStudy (anonymous):

Okay let me write this down

OpenStudy (anonymous):

Okay. NOW I understand

OpenStudy (anonymous):

\[0=x^2-0.20(25)-0.20x\] is now \[0=1^2-0.20(25)-0.20(c)\] Because it says a=1?

OpenStudy (jhannybean):

No the x doesn't just disappear. 1 is a coefficient of x^2 meaning its the integer that is multiplied to the variable x^2. So it would be \(\large 1x^2\) or simply \(\large x^2\)